# Hypothesis testing for community structure in temporal networks using e-values

### v1

@@ 1 Introduction
Despite this extensive work on static networks, to the authors’ knowledge, there are no existing tests for community structure in temporal networks. Perhaps the closest related work comes from network monitoring and surveillance (Woodall et al.,, 2017; Jeske et al.,, 2018). For example, Wilson et al., (2019) study change-point detection in a temporally evolving network using the degree corrected block model (Karrer and Newman,, 2011, henceforth DCBM). Those authors assume that network snapshots are being generated independently and identically distributed until some unknown time $t^{*}$, at which point the data-generating process changes. The goal is to find this change-point time $t^{*}$.
@@ 2.1 Notation
Let $\mathcal{G}=(\mathcal{G}^{(1)},\dots,\mathcal{G}^{(T)})$ be a temporal network where $\mathcal{G}^{(t)}=(\mathcal{V}^{(t)},\mathcal{E}^{(t)})$ is the “snapshot” of the network at time $t$, with node set $\mathcal{V}^{(t)}$ and edge set $\mathcal{E}^{(t)}$, and $T$ is the number of snapshots. At time $t$, we assume that there are $|\mathcal{V}^{(t)}|=n_{t}$ nodes and $|\mathcal{E}^{(t)}|=m_{t}$ edges. Notice that both the number of nodes and edges can change over time, but in this work, we primarily focus on the setting where the number of nodes is constant, i.e., $n_{t}\equiv n$ for all $t$. We also define ${\bf A}=({\bf A}^{(1)},\dots,{\bf A}^{(T)})$ as the adjacency matrix corresponding to $\mathcal{G}$ where $A_{ij}^{(t)}=1$ if nodes $i$ and $j$ have an edge at time $t$, and 0 otherwise. For now, we remain agnostic as to how the network is generated and evolves, but assume that there are no self-loops, i.e., $A_{ii}^{(t)}=0$ for all $i,t$. Lastly, we define $\mathcal{M}_{1}$ as a set of probability distributions and $\mathcal{P}_{0}\subseteq\mathcal{M}_{1}$ as the distributions corresponding to some null hypothesis. We let $\mathbb{P}\in\mathcal{M}_{1}$ represent a single distribution.
@@ 2.3.1 Basics
In words, this condition requires that, for all $\mathbb{P}\in\mathcal{P}_{0}$, $P$ is stochastically no smaller than a uniform random variable on [0,1][0,1][ 0 , 1 ]. On the other hand, an e-value, $E$, is a [0,][0$\infty$][ 0 , ∞ ]-valued random variable such that its expectation under the null hypothesis is at most 1, i.e.,
@@
[…]
@@ 3.1 Proposed test
Since $\mathcal{G}^{(t)}$ is likely dependent on $\mathcal{G}^{(t^{\prime})}$ for $t^{\prime}\neq t$, the corresponding p-values $P_{t}$ and $P_{t^{\prime}}$ will also be dependent. In general, it is quite difficult to combine dependent p-values, especially if this dependence is unknown. To circumvent this issue, we convert each p-value to an e-value since the latter are trivial to combine. To do this, we use a (p-e) calibrator. A calibrator is a decreasing function $g:[0,\infty)\to[0,\infty]$ such that $g(x)=0$ for all $x\in(1,\infty)$ and $g(P)$ is an e-value if $P$ is a p-value. In other words, the calibrator transforms a p-value into an e-value; the reverse transformation is also possible, but we will not need this in the present paper. Some examples of calibrators are given in Shafer et al., (2011) and Vovk and Wang, (2021). Specifically, for any $\kappa\in(0,1)$,
@@
Once the p-values have been converted to e-values, we combine them by taking the arithmetic mean. We showed in Section 2 that not only does this yield a valid e-value, but it can also be shown that this is the best way to combine them with unknown dependence (Vovk and Wang,, 2021). If we knew that the snapshots $\mathcal{G}^{(t)}$ were independent, then we could also multiply the e-values to form a new e-value, but in general we do not expect the network realizations to be independent.
@@ 6 Conclusion
Competing Interests
The authors have no competing interests to declare.
@@
$\tilde{\bf A}=\frac{1}{\sqrt{(n-1)\hat{p}(1-\hat{p})}}({\bf A}-\hat{\bf P}).$
If the null hypothesis is true that the network was generated from an ER model, then Bickel and Sarkar, (2016) show that the largest eigenvalue $\lambda_{1}(\tilde{\bf A})$ (considering negative signs), asymptotically follows the Tracy-Widom distribution with index 1, i.e.,
@@
 $\tilde{\bf A}_{b}\sim\mathsf{ER}(n,\hat{p})$$\tilde{\gamma}_{b}=\frac{1}{\sqrt{(n-1)\hat{p}(1-\hat{p})}}\lambda_{1}(\tilde{\bf A}_{b}-\hat{\bf P})$

### v2

@@ 1 Introduction
Despite this extensive work on static networks, to the authors’ knowledge, there are no existing tests for community structure on temporal networks. Users may be tempted to simply aggregate the temporal network into a static network and then use an existing static hypothesis test for community structure. Not only does this remove any temporal variation, but it has also proven to be a poor approach for other network tasks, e.g., Influence Maximization (Erkol et al.,, 2020; Yanchenko et al.,, 2024). Perhaps the closest related work comes from network monitoring and surveillance (Woodall et al.,, 2017; Jeske et al.,, 2018). For example, Wilson et al., (2019) study change-point detection in a temporally evolving network using the degree corrected block model (Karrer and Newman,, 2011, henceforth DCBM). Those authors assume that network snapshots are being generated independently and identically distributed until some unknown time $t^{*}$, at which point the data-generating process changes. The goal is to find this change-point time $t^{*}$. In another work, Wilson et al., (2017) derive a hypothesis test for community structure in multilayer networks, for which temporal networks could be considered as a special case. Their method identifies densely-connected nodes to compute a significance score which is then compared against the configuration model (Newman,, 2006).
@@ 2.1 Notation
Let $\mathcal{G}=(\mathcal{G}^{(1)},\dots,\mathcal{G}^{(T)})$ be a temporal network where $\mathcal{G}^{(t)}=(\mathcal{V}^{(t)},\mathcal{E}^{(t)})$ is the “snapshot” of the network at time $t$, with node set $\mathcal{V}^{(t)}$ and edge set $\mathcal{E}^{(t)}$, and $T$ is the number of snapshots. At time $t$, we assume that there are $|\mathcal{V}^{(t)}|=n_{t}$ nodes and $|\mathcal{E}^{(t)}|=m_{t}$ edges. In this work, we assume that all edges are undirected, but the ideas can easily be extended to directed networks. Notice that both the number of nodes and edges can change over time, but in this work, we primarily focus on the setting where the number of nodes is constant, i.e., $n_{t}\equiv n$ for all $t$. We also define ${\bf A}=({\bf A}^{(1)},\dots,{\bf A}^{(T)})$ as the adjacency matrix corresponding to $\mathcal{G}$ where $A_{ij}^{(t)}=1$ if nodes $i$ and $j$ have an edge at time $t$, and 0 otherwise. For now, we remain agnostic as to how the network is generated and evolves, but assume that there are no self-loops, i.e., $A_{ii}^{(t)}=0$ for all $i,t$. Lastly, we define $\mathcal{M}_{1}$ as a set of probability distributions over sample space $\Omega$ and $\sigma$-algebra $\mathcal{F}$, and $\mathcal{P}_{0}\subseteq\mathcal{M}_{1}$ as the distributions corresponding to some null hypothesis. We let $\mathbb{P}\in\mathcal{M}_{1}$ represent a single distribution.
@@ 2.3.1 Basics
In words, this condition requires that, for all $\mathbb{P}\in\mathcal{P}_{0}$, $P$ is stochastically no smaller than a uniform random variable on $[0,1]$. On the other hand, an e-value, $E$, is a $0,\infty]-$valued random variable such that its expectation under the null hypothesis is at most 1, i.e.,
@@
where $\mathbb{E}_{\mathbb{P}}$ represents the expectation taken with respect to $\mathbb{P}$. A basic result says that if $E$ is an e-value, then $1/E$ is a p-value. This is easy to see using Markov’s inequality and (2):
@@ 3.1 Proposed test
Since $\mathcal{G}^{(t)}$ is likely dependent on $\mathcal{G}^{(t^{\prime})}$ for $t^{\prime}\neq t$, the corresponding p-values $P_{t}$ and $P_{t^{\prime}}$ will also be dependent. In general, it is quite difficult to combine dependent p-values, especially if this dependence is unknown. To circumvent this issue, we convert each p-value to an e-value since the latter are trivial to combine. To do this, we use a (p-e) calibrator. A calibrator is a non-negative decreasing function $g:[0,\infty)\to[0,\infty]$ with integral at most one such that $g(x)=0$ for all $x\in(1,\infty)$ and $g(P)$ is an e-value if $P$ is a p-value. In other words, the calibrator transforms a p-value into an e-value; the reverse transformation is also possible, but we will not need this in the present paper. Some examples of calibrators are given in Shafer et al., (2011) and Vovk and Wang, (2021). Specifically, for any $\kappa\in(0,1)$,
@@
Once the p-values have been converted to e-values, we combine them by taking the arithmetic mean. We showed in Section 2 that not only does this yield a valid e-value, but it can also be shown that this is the best way to combine them with unknown dependence (Vovk and Wang,, 2021). More generally, for any positive weights $w_{1},\dots,w_{T}$ such that $\sum_{t=1}^{T}w_{t}=1$,
$\tilde{E}_{T}=\sum_{t=1}^{T}w_{t}E_{t}$
is also a valid e-value. Of course, the user would then need to choose the weights, but, e.g., selecting weights such that $w_{1}<w_{2}<\dots<w_{T}$ would place a greater importance on more recent snapshots, which would be reasonable if it was plausible that a community structure was “evolving” in the time index $t$. Finally, if we knew that the snapshots $\mathcal{G}^{(t)}$ were independent, then we could also multiply the e-values to form a new e-value, but in general we do not expect the network realizations to be independent.
@@ 3.2 Properties
Lastly, we stress that the main reason for leveraging an e-value framework is that it easily allows for combining evidence across different snapshots (here, via averaging). If we had used p-values directly, then it would be highly non-trivial to combine results, even if we assume a data-generating model. That being said, by averaging e-values, the proposed approach implicitly tests for an “average” community structure over the temporal evolution of the network. This means, for example, that if the snapshot ordering was shuffled, the proposed approach would still yield the same e-value. We do not consider this a feature or a bug of the method, but rather a reality that users need to be aware of. Indeed, as we discuss further in the Conclusion, defining community structure on a temporal network is quite difficult, and since, to our knowledge, there are no other existing hypothesis tests for this situation, we view this as a sensible first solution. Of course, more explicitly accounting for the evolution in the network could yield a more powerful test. We leave this investigation for future work. Our method does, however, implicitly account for correlation in the network, e.g., if every snapshot was identical (perfect correlation), then the e-value would not be an average and instead would be the value from only a single snapshot.
@@ 6 Conclusion
As another possible direction for future work, our proposed framework can easily be extended to other situations where multiple realizations of the network arise, i.e., multilayer networks. (Kivelä et al.,, 2014). For example, there could be a network where nodes are researchers, and one layer of the network represents collaborations as edges, a different layer uses edges to denote citations, and a third layer corresponds to a friendship network. If the practitioner is interested in testing for the significance of community structure across these different layers, then our approach could easily be applied.
@@
[…]
@@
[…]